OpenStudy (anonymous):

What is the solution of the equation x − 2 = 29?

OpenStudy (unklerhaukus):

add two

OpenStudy (anonymous):

just add 2 on both sides

OpenStudy (anonymous):

x-2+2=29+2 x=31

OpenStudy (anonymous):

Add 2 both the sides and get x value..